# SMaRTT-REPS: Sender-based Marked Rapidly-adapting Trimmed & Timed Transport with Recycled Entropies

### v1

@@ 6. Related Work
Researchers proposed several datacenter CC algorithms, and we discuss in the following how SMaRTT differs from some of the most recently proposed ones. Receiver-based algorithms such as NDP (Handley et al., 2017), EQDS (Olteanu et al., 2022), pHost (Gao et al., 2015), ExpressPass (Zhou et al., 2019), SMSRP (Jiang et al., 2015), and Homa (Montazeri et al., 2018) maintain end-to-end credits that are used to adjust sender flow rates. They can effectively manage incast scenarios by precisely regulating the transmission rates of the different senders. We have shown, however, that they can struggle when dealing with fabric congestion unless they are complemented with a sender-based mechanism. SMaRTT can instead deal with fabric congestion and incast but without requiring pull queues.
On the other hand, sender-based CC algorithms can deal better with fabric congestion, but might not be as responsive when confronted with sudden workload shift (e.g. incast or dynamic job arrivals or departures). Some of those algorithms like DCTCP (Alizadeh et al., 2010b), Hull (Alizadeh et al., 2012), and D2TCP (Vamanan et al., 2012) rely on ECN marking. For example, DCQCN (Zhu et al., 2015) combines ECN and Priority-based Flow Control (PFC) to avoid packet losses and react to congestion quickly. However, PFC is hard to tune and can cause PFC storms (Hoefler et al., 2023; Mittal et al., 2018). Differently from DCQCN, SMaRTT is designed for lossy networks and relies on packet trimming to avoid switches dropping packets, thus avoiding all the issues related to the use of PFC. Other sender-based algorithms Swift (Kumar et al., 2020) and TIMELY (Mittal et al., 2015) detect congestion using delay but, as we show in this paper, they cannot react as fast as ECN-based algorithms.
Last, some works recently demonstrated that having an unfair congestion control can improve performance of jobs running ML training workloads (Rajasekaran et al., 2022). These works are orthogonal to SMaRTT, since they target fairness between different jobs, whereas SMaRTT improves fairness between flows belonging to the same job.
@@ 7. Conclusion
This paper proposes SMaRTT, a sender- and window-based CC algorithm using ECN and delay as congestion signals and packet trimming to detect packet losses. When trimming is not supported SMaRTT falls back to a timeout-based mechanism, which we show adds at most two base RTT of delay. We couple SMaRTT with REPS, a simple and resource-efficient load balancer that significantly improves the FCT, as also shown in our evaluation on a real testbed. REPS can drastically improve the load balancing performance compared to traditional per-flow ECMP or oblivious packet spraying at minimal complexity cost. We also show that REPS can effectively handle faults and network asymmetries using a real network. We evaluate SMaRTT on several workloads, showing that it constantly either outperforms or matches the performance of EQDS, Swift, BBR, and MPRDMA by up to 50%.

### v3

@@ 6. Related Work
Researchers proposed several datacenter CC algorithms, and we discuss in the following how FASTFLOW differs from some of the most recently proposed ones. Receiver-based algorithms such as NDP (Handley et al., 2017), EQDS (Olteanu et al., 2022), pHost (Gao et al., 2015), ExpressPass (Zhou et al., 2019), SMSRP (Jiang et al., 2015), and Homa (Montazeri et al., 2018) maintain end-to-end credits that are used to adjust sender flow rates. They can effectively manage incast scenarios by precisely regulating the transmission rates of the different senders. We have shown, however, that they can struggle when dealing with fabric congestion unless they are complemented with a sender-based mechanism. FASTFLOW can instead deal with fabric congestion and incast but without requiring pull queues.
On the other hand, sender-based CC algorithms can deal better with fabric congestion, but might not be as responsive when confronted with sudden workload shift (e.g. incast or dynamic job arrivals or departures). Some of those algorithms like DCTCP (Alizadeh et al., 2010b), Hull (Alizadeh et al., 2012), and D2TCP (Vamanan et al., 2012) rely on ECN marking. For example, DCQCN (Zhu et al., 2015) combines ECN and Priority-based Flow Control (PFC) to avoid packet losses and react to congestion quickly. However, PFC is hard to tune and can cause PFC storms (Hoefler et al., 2023; Mittal et al., 2018). Differently from DCQCN, FASTFLOW is designed for lossy networks and relies on packet trimming to avoid switches dropping packets, thus avoiding all the issues related to the use of PFC. Other sender-based algorithms Swift (Kumar et al., 2020) and TIMELY (Mittal et al., 2015) detect congestion using delay but, as we show in this paper, they cannot react as fast as ECN-based algorithms.
Last, some works recently demonstrated that having an unfair congestion control can improve performance of jobs running ML training workloads (Rajasekaran et al., 2022). These works are orthogonal to FASTFLOW, since they target fairness between different jobs, whereas FASTFLOW improves fairness between flows belonging to the same job.
@@ 7. Conclusion
This paper proposes FASTFLOW, a sender- and window-based CC algorithm using ECN and delay as congestion signals and packet trimming to detect packet losses. When trimming is not supported FASTFLOW falls back to a timeout-based mechanism, which we show adds at most two base RTT of delay. We evaluate FASTFLOW on several workloads, showing that it constantly either outperforms or matches the performance of EQDS, Swift, BBR, and MPRDMA by up to 50%. Eventually, we show how FASTFLOW could complement receiver-based CC algorithms such as EQDS.